### v2

# Multiscale Stochastic Gradient Descent: Efficiently Training Convolutional Neural Networks

Abstract
Stochastic Gradient Descent (SGD) is the foundation of modern deep learning optimization but becomes increasingly inefficient when training convolutional neural networks (CNNs) on high-resolution data. This paper introduces Multiscale Stochastic Gradient Descent (Multiscale-SGD), a novel optimization approach that exploits coarse-to-fine training strategies to estimate the gradient at a fraction of the cost, improving the computational efficiency of SGD type methods while preserving model accuracy. We derive theoretical criteria for Multiscale-SGD to be effective, and show that while standard convolutions can be used, they can be suboptimal for noisy data. This leads us to introduce a new class of learnable, scale-independent Mesh-Free Convolutions (MFCs) that ensure consistent gradient behavior across resolutions, making them well-suited for multiscale training. Through extensive empirical validation, we demonstrate that in practice, (i) our Multiscale-SGD approach can be used to train various architectures for a variety of tasks, and (ii) when the noise is not significant, standard convolutions benefit from our multiscale training framework. Our results establish a new paradigm for the efficient training of deep networks, enabling practical scalability in high-resolution and multiscale learning tasks.
Machine Learning, ICML
Niloufar Zakariaei1∗, Shadab Ahamed1∗, Eldad Haber1, Moshe Eliasof2
1
University of British Columbia, Vancouver, Canada
2University of Cambridge, Cambridge, United Kingdom
nilouzk@student.ubc.ca, shadab@phas.ubc.ca, ehaber@eoas.ubc.ca, me532@cam.ac.uk
* Equal contribution
1 Introduction
In this work, we consider the task of learning a functional $y({\bf x})=\phi(u({\bf x}))$, where ${\bf x}$ is a position (in 2D ${\bf x}=({\bf x}_{1},{\bf x}_{2})$ and in 3D ${\bf x}=({\bf x}_{1},{\bf x}_{2},{\bf x}_{3})$), $u({\bf x})\in{\cal U}$ and $y({\bf x})\in{\cal Y}$ are families of functions. To this end, we assume to have discrete samples from ${\cal U}$ and ${\cal Y}$, $({\bf u}_{i}^{h}=u_{i}({\bf x}_{h}),{\bf y}^{h}_{i}=\phi(u_{i}({\bf x}_{h}))),\ \ i=1,\ldots,M$ associated with some resolution $h$. A common approach to learning the function is to parameterize the problem, typically by a deep network, and replace $\phi$ with a function $f(\cdot,\cdot)$ that accepts the vector ${\bf u}^{h}$ and learnable parameters ${\bm{\theta}}$ which leads to the problem of estimating ${\bm{\theta}}$ such that, ${\bf y}^{h}_{i}\approx f({\bf u}^{h}_{i},{\bm{\theta}})$ for $i=1,\ldots,M$. To evaluate ${\bm{\theta}}$, the following stochastic optimization problem is solved:
$\min_{{\bm{\theta}}}\ {\mathbb{E}}_{{\bf u}^{h},{\bf y}^{h}}{\ell}\left(f({\bf u}^{h},{\bm{\theta}}),{\bf y}^{h}\right).$ (1)
Here, $\ell(\cdot,\cdot)$ is a loss function, typically the mean square error. Standard approaches use variants of stochastic gradient descent (SGD) to estimate the loss and its gradient for different samples of $({\bf u}^{h},{\bf y}^{h})$. In deep learning with convolutional neural networks, the parameter ${\bm{\theta}}$ (the convolutional weights) has identical dimensions, independent of the resolution. Although SGD is widely used, its computational cost can become prohibitively high as the mesh-size $h$ decreases, especially when evaluating the function $f$ on a fine mesh for many samples ${\bf u}_{i}^{h}$. This challenge is worsened if the initial guess ${\bm{\theta}}$ is far from optimal, requiring many costly iterations, for large data sizes $M$. One way to avoid large meshes is to use small crops of the high resolution image, however, this can degrade performance, especially when a large receptive field is required for learning (Araujo et al., 2019).
Background and related work. Computational cost reduction can be achieved by leveraging different resolutions, a concept foundational to multigrid and multiscale methods. These methods have a long history of solving partial differential equations and optimization problems (Trottenberg et al., 2001; Briggs et al., 2000; Nash, 2000). Techniques like multigrid (Trottenberg et al., 2001) and algorithms such as MGopt (Nash, 2000; Borzi, 2005) and Multilevel Monte Carlo (Giles, 2015, 2008; Van Barel & Vandewalle, 2019) are widely used for optimization and differential equations.
In deep learning, multiscale or pyramidal approaches have been used in image processing tasks such as object detection, segmentation, and recognition, where analyzing multiple resolutions is key (Scott & Mjolsness, 2019; Chada et al., 2022), reviewed in Elizar et al. (2022). Recent methods improve standard CNNs for multiscale computations by introducing specialized architectures and training methods. For instance, He & Xu (2019) uses multigrid methods in CNNs to boost efficiency and capture multiscale features, while Eliasof et al. (2023b) focuses on multiscale channel space learning, and van Betteray et al. (2023) unifies both. Li et al. (2020b) introduced the Fourier Neural Operator, enabling mesh-independent computations, and Wavelet-NNs were explored to capture multiscale features via wavelets (Fujieda et al., 2018; Finder et al., 2022; Eliasof et al., 2023a).
While often overlooked, it is important to note that these approaches, can be divided into two families of approaches that leverage multiscale concepts. The first is to learn parameters for each scale, and a separate set of parameters that mix scales, as in UNet (Ronneberger et al., 2015). The second, called multiscale training, enables the approximation of fine-scale parameters using coarse-scale samples (Haber et al., 2017; Wang et al., 2020; Ding et al., 2020; Ganapathy & Liew, 2008). The second approach aims to gain computational efficiency, as it approximates fine mesh parameters using coarse meshes, and it can be coupled with the first approach, and in particular with UNets.
Our approach. This work falls into the second category of multiscale training. We study multiscale algorithms that use coarse meshes to approximate high-resolution quantities, particularly the gradients of network parameters. Computing gradients on coarse grids is significantly cheaper than on fine grids, as noted in (Shi & Cornish, 2021). However, for efficient multiscale training, parameters on coarse and fine meshes must have “similar meaning," implying that both the loss and gradient on coarse meshes should approximate those on fine meshes. Specifically, the loss and gradient with respect to parameters should converge to a finite value as $h\rightarrow 0$. In this work, we show that standard CNN gradients may not converge as the mesh size $h$ approaches 00, suggesting CNNs under-utilize multiscale training. This motivates our development of mesh-free convolution kernels, whose values and gradients converge as $h\rightarrow 0$. Our approach builds on Differential-Operator theory (Wong, 2014) to create a family of learnable, mesh-independent convolutions for multiscale learning, resembling Fourier Neural Operators (FNO) (Li et al., 2020a) but with enhanced expressiblity.
Our main contributions are: 1. Propose a new multiscale training algorithm, Multiscale SGD. 2. Analyze the limitations of standard CNNs within a multiscale framework. 3. Introduce a family of mesh-independent CNNs inspired by differential operators. 4. Validate our approach on benchmark tasks, showcasing enhanced efficiency and scalability.
Figure 1: Illustration of our Multiscale-SGD, introduced in Section 2.
2 Multiscale Stochastic Gradient Descent
We now present the standard approach for training CNNs, identify its major computational bottleneck, and propose a novel solution called Multiscale Stochastic Gradient Descent (Multiscale-SGD).
Standard training of neural networks. Suppose we use a gradient descent-based method to train a CNN, with input resolution [^1] In this paper, we define resolution $h$ as the pixel size on a 2D uniform meshgrid, where a smaller $h$ indicates a higher resolution. For simplicity, we assume the same $h$ across all dimensions, though different resolutions can be assigned per dimension. $h$ with trainable parameters ${\bm{\theta}}$. At $k^{\text{th}}$ iteration, we have
${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}-\mu_{k}{\mathbb{E}}_{{\bf u}^{h},{\bf y}^{h}}\left[{\bf g}({\bf u}^{h},{\bf y}^{h},{\bm{\theta}}_{k})\right],$ (2)
where the gradient of the loss $\ell$ with respect to ${\bm{\theta}}$ is ${\bf g}({\bf u}^{h},{\bf y}^{h},{\bm{\theta}})={\bm{\nabla}\,}\ell\left(f({\bf u}^{h},{\bm{\theta}}_{k}),{\bf y}^{h}\right)$. Evaluating the expected value of the gradient can be highly expensive. To understand why, consider the estimation of the gradient obtained via the average over ${\bf g}$ with a batch of $N$ samples,
${\mathbb{E}}_{{\bf u}^{h},{\bf y}^{h}}\,\left[{\bf g}({\bf u}^{h},{\bf y}^{h},{\bm{\theta}})\right]\approx{\frac{1}{N}}\sum_{i}{\bf g}({\bf u}_{i}^{h},{\bf y}_{i}^{h},{\bm{\theta}}_{k}).$ (3)
This approximation results in an error in the gradient. Under some mild assumptions on the sampling of the gradient value (see Johansen et al. (2010)), the error can be bounded by,
$\left\|{\mathbb{E}}\,\left[{\bf g}({\bf u}^{h},{\bf y}^{h},{\bm{\theta}}_{k})\right]-{\frac{1}{N}}\sum_{i}{\bf g}({\bf u}_{i}^{h},{\bf y}_{i}^{h},{\bm{\theta}}^{k})\right\|_{2}\leq{\frac{C}{\sqrt{N}}},$ (4)
where $C$ is some constant. Clearly, obtaining an accurate evaluation of the gradient (that is, with low variance) requires sampling ${\bf g}$ across many data points with sufficiently high resolution $h$. This tradeoff between the sample size $N$ and the accuracy of the gradient estimation, is the costly part of training a deep network on high-resolution data. To alleviate the problem, it is common to use large batches, effectively enlarging the sample size. It is also possible to use variance reduction techniques (Anschel et al., 2017; Chen et al., 2017; Alain et al., 2015). Nonetheless, for high-resolution images, or 3D inputs, the large memory requirement limits the size of the batch. However, a small batch size can result in noisy, highly inaccurate gradients, and slow convergence (Shapiro et al., 2009).
Figure 2: Illustration of the Full Multiscale training using the Full Multiscale Stochastic Gradient Descent (MSGD) algorithm described in Section 4.
2.1 Efficient training with Multiscale Stochastic Gradient Descent
To reduce the cost of the computation of the gradient, we use a classical trick proposed in the context of Multilevel Monte Carlo methods (Giles, 2015). To this end, let $h=h_{1}<h_{2}<...<h_{L}$ be a sequence of mesh step sizes, in which the functions $u$ and $y$ are discretized with. We can easily sample (or coarsen) $u$ and $y$ to some mesh $h_{j},1\leq j\leq L$. We consider the following identity, based on the telescopic sum and the linearity of the expectation,
${\mathbb{E}}\left[{\bf g}^{h_{1}}({\bm{\theta}})\right]$ $={\mathbb{E}}\left[{\bf g}^{h_{L}}({\bm{\theta}})\right]+{\mathbb{E}}\left[{\bf g}^{h_{L-1}}({\bm{\theta}})-{\bf g}^{h_{L}}({\bm{\theta}})\right]$
$+\ldots\ +{\mathbb{E}}\left[{\bf g}^{h_{1}}({\bm{\theta}})-{\bf g}^{h_{2}}({\bm{\theta}})\right],$ (5)
where for shorthand we define the gradient of ${\bm{\theta}}$ with resolution $h_{j}$ by ${\bf g}^{h_{j}}({\bm{\theta}})={\bf g}({\bf u}^{h_{j}},{\bf y}^{h_{j}},{\bm{\theta}})$. The core idea of our Multiscale Stochastic Gradient Descent (Multiscale-SGD) approach, is that the expected value of each term in the telescopic sum is approximated using a different batch of data with a different batch size. This concept is demonstrated in Figure 1 and it can be summarized as,
${\mathbb{E}}\left[{\bf g}^{h_{1}}({\bm{\theta}})\right]$ $\approx{\frac{1}{N_{L}}}\sum_{i}{\bf g}^{h_{L}}_{i}({\bm{\theta}})$
$+{\frac{1}{N_{L-1}}}\sum_{i}\left({\bf g}^{h_{L-1}}_{i}({\bm{\theta}})-{\bf g}^{h_{L}}_{i}({\bm{\theta}})\right)$
$+\ldots+{\frac{1}{N_{1}}}\sum_{i}\left({\bf g}^{h_{1}}_{i}({\bm{\theta}})-{\bf g}^{h_{2}}_{i}({\bm{\theta}})\right).\vspace{-0.5cm}$ (6)
To understand why this concept is beneficial, we analyze the error obtained by sampling each term in Section 2.1. Evaluating the first term in the sum requires evaluating the function ${\bf g}$ on the coarsest mesh (i.e., lowest resolution) using downsampled inputs. Therefore, it can be efficiently computed, while utilizing a large batch size, $N_{L}$. Thus, following Equation 4, the approximation error of the first term in the Section 2.1 can be bounded by,
$\left\|{\mathbb{E}}\left[{\bf g}^{h_{L}}({\bm{\theta}})\right]-{\frac{1}{N_{L}}}\sum_{i}{\bf g}^{h_{L}}_{i}({\bm{\theta}})\right\|_{2}\leq{\frac{C}{\sqrt{N_{L}}}}.$ (7)
Following this step, we need to evaluate the terms of the form,
${\bf r}_{j}={\mathbb{E}}\left[{\bf g}^{h_{j-1}}({\bm{\theta}})-{\bf g}^{h_{j}}({\bm{\theta}})\right],$ (8)
Similarly, this step can be computed by resampling, with a batch size $N_{j}$,
$\hat{{\bf r}}_{j}={\frac{1}{N_{j-1}}}\sum_{i}\left({\bf g}^{h_{j-1}}_{i}({\bm{\theta}})-{\bf g}^{h_{j}}_{i}({\bm{\theta}})\right),$ (9)
for $j=2,\ldots,L-1$. The key question is: what is the error in approximating ${\bf r}_{j}$ by the finite sample estimate $\hat{{\bf r}}_{j}$? Previously, we focused on errors due to sample size. However, note that the exact term ${\bf r}_{j}$ is computed by evaluating ${\bf g}$ on two resolutions of the same samples and subtracting the results.
If the evaluation of ${\bf g}$ on different resolutions yields similar results, then ${\bf g}$ computed on mesh with step size $h_{j}$ can be utilized to approximate the gradient ${\bf g}$ on mesh a mesh with finer resolution $h_{j-1}$, making the approximation error $\hat{{\bf r}}_{j}$ small. Furthermore, assume that
$\|{\bf g}_{i}^{h_{j-1}}({\bm{\theta}})-{\bf g}_{i}^{h_{j}}({\bm{\theta}})\|\leq Bh_{j}^{p}\quad p>0,$ (10)
for some constant $B>0$ and $p>0$, both independent of the pixel-size $h_{j}$. Then, we can bound the error of approximating ${\bf r}_{j}$ by $\hat{{\bf r}}_{j}$ as,
$\|{\bf r}_{j}-\hat{{\bf r}}_{j}\|\leq BC{\frac{h_{j}^{p}}{\sqrt{N_{j}}}}.$ (11)
Note that, under the assumption that Equation 10 holds, the gradient approximation error between different resolutions decreases as the resolution increases (i.e., $h\rightarrow 0)$. Indeed, the sum of the gradient approximation error between subsequent resolutions (where each is defined in Equation 8), where the approximation is obtained from the telescopic sum in Section 2.1, can be bounded by,
$e=C\left({\frac{1}{\sqrt{N_{L}}}}+B\sum_{j=1}^{L-1}{\frac{h_{j}^{p}}{\sqrt{N_{j}}}}\right).$ (12)
Let us look at an exemplary case, where $p=1$ and $h_{j}=2h_{j-1}$, i.e., the resolution on each dimension increases by a factor of 2 between input representations. In this case, the sampling error on the coarsest mesh contributes $N_{L}^{-1/2}$. It then follows that, it is also possible to have the same order of error by choosing $N_{L-1}=N_{L}/4$. That is, to obtain the same order of error at subsequent levels, only $1/4$ of the samples are required at the coarser grid compared to the finer one.
Following our Multiscale-SGD approach in Section 2.1, the sample size needed on the finest mesh is reduced by a factor of $4^{L}$ from the original $N_{L}$ while maintaining the same error order, leading to significant computational savings. For instance, Table 3 shows that using 4 mesh resolution levels achieves orders of magnitude in computational savings compared to a single high-resolution gradient approximation. Beyond these savings, Multiscale-SGD is easy to implement. It simply requires computing the loss at different input scales and batches, which can be done in parallel. Since gradients are linear, the loss gradient naturally yields Multiscale-SGD. The full algorithm is outlined in Algorithm 1.
Algorithm 1 Multiscale Stochastic Gradient Descent (Multiscale-SGD)
  Set batch size to $N_{L}$ and sample $N_{L}$ samples of ${\bf u}^{h_{1}}$ and ${\bf y}^{h_{1}}$
  Pool ${\bf u}^{h_{L}}={\bf R}^{h_{1}}_{h_{L}}{\bf u}^{h_{1}}$,  ${\bf y}^{h_{L}}={\bf R}^{h_{1}}_{h_{L}}{\bf y}^{h_{1}}$
  Set $loss=\ell({\bf u}^{h_{L}},{\bf y}^{h_{L}},{\bm{\theta}})$
  for $j=1,\dots,L$ in parallel do
     Set batch size to $N_{j}$ and sample $N_{j}$ samples of ${\bf u}^{h_{1}}$ and ${\bf y}^{h_{1}}$
     Pool ${\bf u}^{h_{j}}={\bf R}^{h_{1}}_{h_{j}}{\bf u}^{h_{1}}$,  ${\bf y}^{h_{j}}={\bf R}^{h_{1}}_{h_{j}}{\bf y}^{h_{1}}$
     Compute losses:  $\ell({\bf u}^{h_{j}},{\bf y}^{h_{j}},{\bm{\theta}})$ and $\ell({\bf u}^{h_{j-1}},{\bf y}^{h_{j-1}},{\bm{\theta}})$
     $loss\leftarrow loss-\ell({\bf u}^{h_{j}},{\bf y}^{h_{j}},{\bm{\theta}})+\ell({\bf u}^{h_{j-1}},{\bf y}^{h_{j-1}},{\bm{\theta}})$
  end for
2.2 Multiscale analysis of convolutional neural networks
We now analyze under which conditions Multiscale-SGD can be applied for neural network optimization without compromising on its efficiency.
[…]
Lemma 2.1 (Convergence of standard convolution kernels).
[…]
$\|{\bf g}^{2h}-{\bf g}^{h}\|\sim{\cal O}(h).$
As shown in the proof (see Appendix B), the convergence of the gradients depends on the amount of noise in the data. However, as we now show in Example 1, for noisy problems, this method may not always yield the desired outcome.
Example 1.
Assume that $f$ is a 1D convolution, that is $f({\bf u}^{h},{\bm{\theta}})={\bf u}^{h}\star{\bm{\theta}},$ and a linear model
$\ell_{h}({\bm{\theta}})={\frac{1}{n}}({\bf u}^{h}\star{\bm{\theta}})^{\top}{\bf y}^{h},$ (13)
where ${\bf y}^{h}$ is the discretization of some function $y$ on the fine mesh, $h$. We use a 1D convolution with kernel size 3, whose trainable weight vector is ${\bm{\theta}}\in\mathbb{R}^{3\times 1}$ and compute the gradient of the function $y$ on different meshes. The loss function on the $i$-th mesh can be written as
$\ell_{2^{i}h}({\bm{\theta}})={\frac{1}{n_{i}}}({\bf R}_{h}^{2^{i}h}{\bf u}^{h}\star{\bm{\theta}})^{\top}({\bf R}_{h}^{2^{i}h}{\bf y}^{h}),$ (14)
where ${\bf R}_{h}^{2^{i}h}$ is a linear interpolation operator that takes the signal from mesh size $h$ to mesh size $2^{i}h$. We compute the difference between the gradient of the function on each level and its norm, $\delta g=\|{\bm{\nabla}\,}_{{\bm{\theta}}}\ell_{2^{i}h}-{\bm{\nabla}\,}_{{\bm{\theta}}}\ell_{2^{i+1}h}\|^{2}$.
We conduct two experiments: in the first, we use smooth inputs, and in the second, we use smooth inputs with added Gaussian noise $\mathcal{N}(0,1)$. We compute the values of $\delta g$, and report it in Table 1. The experiment demonstrates that for noisy inputs, the gradient with respect to the convolution parameters grows in tandem with the mesh. Thus, using convolution in its standard form will not be useful in reducing the error using the telescopic sum in Section 2.1 as described above.
Table 1: The discrepancy between two consecutive gradients at different resolutions for the loss in Equation 13 for standard convolutions. As the mesh resolution increases, the discrepancy reduces for smooth data, while it increases for noisy data.
Computed loss on different resolutions (×$\times 10^{-3}$)
Resolution 1/1024 1/512 1/256 1/128 1/64 1/32
Noisy input 9.3 7.1 4.9 2.8 2.0 1.7
Smooth input 0.12 0.25 0.51 1.0 1.9 3.7
The analysis suggests that for smooth signals, standard CNNs can be integrated into Multilevel Monte-Carlo methods, which form the basis of our Multiscale-SGD. However, in many cases, inputs may be noisy or contain high frequencies, leading to large or unbounded gradients across mesh resolutions. This explains the lack of convergence of standard CNNs in Example 1. High-frequency inputs can hinder learning in multiscale frameworks like Multiscale-SGD. To address this, we propose a new convolution type in Section 3 that theoretically alleviates this issue when dealing with very noisy data.
3 Multiscale Training with Mesh-Free Convolutions
In Section 2.2, we analyzed standard convolutional kernels when coupled with our Multiscale-SGD training approach. As we have shown, standard convolutional kernels can scale poorly in the presence of noisy data. To address this limitation, we now propose an alternative to standard convolutions through the notion of ’mesh-free’ or weakly mesh-dependent convolution kernels.
3.1 Mesh-Free Convolutions
We introduce Mesh-Free Convolutions (MFCs), which define convolutions in the functional space independently of any mesh, then discretize them on a given mesh obtaining consistency. Unlike standard convolution kernels tied to specific meshes, MFCs converge to a finite value as the mesh refines, overcoming the limitations of standard kernels discussed in Section 2.2. One way to achieve this goal is to define the convolution in functional space directly by an integral,
[…]
where the kernel function ${\cal K}_{{\bm{\xi}}}$ is, parametrized by learnable parameters ${\bm{\xi}}$. This general convolution can be defined on any mesh and is basically mesh-free. The question remains: how to parametrize the kernel? Clearly, there are many options for this choice. A too-restrictive approach leads to limited expressivity of the convolution. Our goal and intuition is to generate kernels that are commonly obtained by trained CNNs. Such filters are rarely random but contain some structure (Goodfellow et al., 2016). To this end, we use the interplay between the solution of partial differential equations (PDEs) and convolutions (Evans, 1998). In particular, the solution of parabolic PDEs (that is, the heat equation) is a kernel that exhibits the behavior that is often required from CNN filters.
Let $u$ be input functions (that is, $u$ can have more than a single channel, or so-called vector function) and let $v={\cal C}({\bm{\xi}})u$ be the application of the mesh-free convolution is parameterized by ${\bm{\xi}}$. We obtain a mesh-independent convolution by composing two processes. The first is a trainable matrix ${\bf R}\in\mathbb{R}^{d\times c}$ which is a $1\times 1$ convolution, that embeds $u$ into a subspace ${\tilde{v}}$, yielding the initial condition of the PDE defined in Section 3.1. The second is the solution of the parabolic PDE of the form below, which maps the initial condition $\tilde{v}(0,{\bf x})$ into $\tilde{v}(\tau,{\bf x})$
$\tilde{v}_{t}={\cal L}({\bm{\alpha}})\tilde{v},\quad\tilde{v}(t=0)={\bf R}u,\quad t\in[0,\tau],\quad\quad$
$\quad v=(\gamma_{0}+[\gamma_{x},\gamma_{y}]^{\top}{\bm{\nabla}\,})\tilde{v}({\bf x},\tau).$ (16)
Finally, the output $v({\bf x})$ is obtained by taking a directional derivative of $\tilde{v}(\tau,{\bf x})$ in a direction ${\bm{\gamma}}$. All the processes above are encompassed by the mesh-free convolution operator ${\cal C}({\bm{\xi}})$.
The operator ${\cal L}({\bm{\alpha}})={\rm diag}\left({\cal L}({\bm{\alpha}}_{1}),\ldots,{\cal L}({\bm{\alpha}}_{d})\right)$ is an Elliptic Differential Operator, that we define as
${\cal L}=\alpha_{xx}{\frac{\partial^{2}v}{\partial x^{2}}}+2\alpha_{xy}{\frac{\partial^{2}v}{\partial xy}}+\alpha_{yy}{\frac{\partial^{2}v}{\partial y^{2}}}.$ (17)
The trainable parameters are ${\bm{\xi}}=[\tau,{\bm{\alpha}},{\bm{\gamma}}]\in\mathbb{R}^{7}$, where $\tau$ is a scalar, ${\bm{\gamma}}=[\gamma_{0},\gamma_{x},\gamma_{y}]$, and ${\bm{\alpha}}=[\alpha_{xx},\alpha_{xy},\alpha_{yy}]$. We enforce $\alpha_{xx},\alpha_{yy}>0$ and $\alpha_{xx}\alpha_{yy}-\alpha_{xy}^{2}>0$, ensuring ${\cal L}$ is elliptic and invertible (Trottenberg et al., 2001).
Integrating the PDE over time $[0,\tau]$ and differentiating in the direction $[\gamma_{x},\gamma_{y}]$ effectively applies the convolution to the input $u$. The parameter $\tau$ acts as a scaling factor: increasing it widens and smooths the kernel, while as $\tau\to 0$, the kernel becomes more localized. To visualize our convolution, we sample random parameters ${\bm{\xi}}$ and plot the resulting MFC filters in Figure 3. By varying ${\bm{\xi}}$, we generate a range of convolutional kernels, from local to global, with different orientations. Here, $\alpha$ controls orientation, $\gamma$ sets the lobe direction, and $\tau$ determines the kernel width. The figure demonstrates that our choice of convolutions yields trainable filters to the ones that are typically obtained by CNN (Goodfellow et al., 2016). These filters can change their field of view by changing $\tau$, change their orientation by changing ${\bm{\xi}}$, and change their side lobes’ strength and direction by changing ${\bm{\gamma}}$.
Figure 3: An illustration of Mesh-Free Convolutions (MFCs) to showcase the effect of parameter variations. First row: The parameters are set as $\alpha_{xx}=1$, $\alpha_{yy}=3$, while $\alpha_{xy}$ varies from 0 to 1, and ${\bm{\gamma}}_{x}$ changes from 0 to 1. Second row: The parameters are set as ${\bm{\gamma}}_{x}={\bm{\gamma}}_{y}=0$, ${\bm{\gamma}}_{0}=1$, with $\alpha_{xy}$ varying from 0 to 1, and $t$ changing from 0 to 1. These visualizations highlight the flexibility and expressiveness of MFCs in adapting to different parameter configurations.
Computing MFCs on a discretized mesh. Our MFCs are defined by the solution of Section 3.1 and are continuous functions. Here, we discuss their computation on a discretized mesh for application to images. For a rectangular domain $\Omega$, we use periodic boundary conditions and the Fast Fourier Transform (FFT) to compute the convolution. Let ${\bf u}^{h}$ be a discrete tensor. First, we compute its Fourier transform, $\widehat{\bf u}^{h}={{\bf F}}{\bf u}^{h}$, and use a frequency mesh to obtain the 2D frequency grid ${\bf k}_{x}$ and ${\bf k}_{y}$. The discrete symbol ${\cal L}({\bf k}_{x},{\bf k}_{y})$ is a 3D tensor operating on each channel, similar to standard 2D filters. We then compute a point-wise product and use the inverse FFT to return to image space. Interestingly, given some fixed mesh of size $h$ and a regular 2D convolution ${\bf C}$ it is always possible to build an MFC, ${\bf C}_{{\bm{\xi}}}$, such that they are equivalent when evaluated on that mesh. The proof is given in Appendix B, and it implies that MFCs have at least the same expressiveness as standard convolutions.
Computational complexity. MFC first applies a standard pointwise operation to expand a $d$-channel tensor into a $c$-channel tensor. The main cost comes from the spatial operations on the $c$ channels, particularly the FFT, which has a cost of $n\log(n)$, where $n$ is the number of pixels. With $c$ channels, the total cost becomes ${\cal O}(cn\log(n))$. As other operations in the network scale linearly with $n$, the overall cost is dominated by the FFT.
3.2 Fourier analysis of Mesh-Free Convolutions
We now analyze the ability of MFCs to produce consistent gradients across scales, enabling effective use of Multiscale-SGD from Section 2. Unlike standard convolutions, MFCs are defined as a continuous function, not tied to a specific mesh. This allows us to compute gradients directly with respect to the function, equivalent to computing gradients at the finest resolution, as $h\rightarrow 0$.
Following the analysis path shown in Section 2.2, consider the linear loss obtained by these MFCs when discretizing the functions $u$ and $y$. In the case of functions, such as with our MFCs, the inner product in Equation 13 is replaced by an integral,
[…]
We study the behavior of our MFC, denoted by ${\cal C}({\bm{\xi}})$, on a function $u(x)$, by computing its symbol. Because the heat equation is linear, the symbol of ${\cal C}({\bm{\xi}})$ is given by (see (Trottenberg et al., 2001)),
${\cal L}(k_{x},$ $k_{y})=\left(\gamma_{0}+ik_{x}\gamma_{x}+ik_{y}\gamma_{y}\right)\cdot$
$\exp\left(-\tau(\alpha_{xx}k_{x}^{2}+2\alpha_{xy}k_{x}k_{y}+\alpha_{yy}k_{y}^{2})\right).$ (19)
We now provide a Lemma that characterizes the behavior of our MFCs and, in particular, shows that unlike with traditional convolutional operations, our MFCs yield consistent gradient approximations independently of the mesh resolution – thereby satisfying the assumptions of our Multiscale-SGD training approach. The proof is provided in Appendix B.
Lemma 3.1 (MFCs yield consistent gradients independently of mesh resolution.).
Let the loss be defined as Equation 18 and let $u(x)$ and $y(x)$ be any two integrable functions. Then, $-\infty<\ell^{*}<\infty$, that is, the loss of any two integrable functions is a finite number. Furthermore, let ${\bm{\nabla}\,}_{{\bm{\xi}}}\ell^{*}$ be the gradient of the loss with respect to its parameters. Then we have that $-\infty<{\bm{\nabla}\,}_{{\bm{\xi}}}\ell^{*}<\infty$, that is, its gradient is a vector with finite values.
The property described in Lemma 3.1 carries forward to any discrete space, i.e., chosen mesh. Namely, by discretizing our MFCs on a uniform mesh with mesh-size $h$ we obtain its discrete analog:
$\ell({\bm{\xi}})={\frac{1}{n}}({\bf y}^{h})^{\top}({\cal C}({\bm{\xi}}){\bf u}^{h}).$ (20)
Furthermore, using the discrete Fourier Transform ${\bf F}$, we obtain that
$\ell^{h}({\bm{\xi}})={\frac{1}{n}}({\bf F}{\bf y}^{h})^{\top}{\cal L}(k_{x},k_{y})({\bf F}{\bf u}^{h})$
$={\frac{1}{n}}\sum_{k_{x},k_{y}}{\cal L}(k_{x},k_{y})\odot\widehat{\bf y}^{h}(k_{x},k_{y})\odot\widehat{\bf u}^{h}(k_{x},k_{y}).$ (21)
As both $\widehat{\bf y}^{h}$ and $\widehat{\bf u}^{h}$ are discrete Fourier transforms of the functions $y$ and $u$, they converge to their continuous counterparts as $h\rightarrow 0$. Thus, $\ell^{h}$ converges to $\ell^{*}$ as $h\rightarrow 0$, leading to a key result: unlike standard convolutions, the gradients of our MFCs converge to a finite value as the mesh refines, making them suitable for multiscale techniques like Multiscale-SGD, summarized in Lemma 3.2.
Lemma 3.2 (MFCs gradients converge as the mesh resolution refines).
Let $\ell^{h}({\bm{\xi}})$ be defined by Equation 20. At the limit,
$\lim_{h\rightarrow 0}\ell=\ell^{*}\quad{\rm and}\quad\lim_{h\rightarrow 0}{\bm{\nabla}\,}_{{\bm{\xi}}}\ell={\bm{\nabla}\,}_{{\bm{\xi}}}\ell^{*}.$
To demonstrate this property of the mesh-free convolution, we repeat Example 1, but this time with the mesh-free convolutions. The results are reported in Table 2.
Table 2: The discrepancy between two consecutive gradients at different resolutions for the loss in Equation 13 for mesh-free convolutions (MFCs). The discrepancy is reduced for both noisy and smooth data.
Computed loss on different resolutions (×$\times 10^{-3}$)
Resolution 1/1024 1/512 1/256 1/128 1/64 1/32
Noisy input 0.16 0.32 0.61 1.3 2.9 5.9
Smooth input 0.11 0.21 0.42 0.78 1.5 3.1
4 The Full-Multiscale Algorithm
To leverage a multiscale framework, a common method is to first solve on a coarse mesh and interpolate it to a finer mesh, a process called mesh homotopy (Haber et al., 2007). When training a neural network, optimizing the weights ${\bm{\theta}}$ on coarse meshes requires fewer iterations to refine on fine meshes (Borzì & Schulz, 2012). Using Multiscale-SGD (Algorithm 1), the number of function evaluations on fine meshes is minimized. This resolution-dependent approach is summarized in Full-Multiscale-SGD (Borzì & Schulz, 2012) (Algorithm 2) and illustrated in Figure 2.
Algorithm 2 Full-Multiscale-SGD
  Randomly initialize the trainable parameters ${\bm{\theta}}_{*}^{H}$.
  for $j=1,...,L$ do
     Set mesh size to $h_{j}=2^{L-j}h$ and ${\bm{\theta}}_{0}^{h_{j}}={\bm{\theta}}_{*}^{H}$.
     Solve the optimization problem on mesh $h_{j}$ using SGD and Algorithm 1 for ${\bm{\theta}}_{*}^{h_{j}}$.
     Set ${\bm{\theta}}_{*}^{H}\leftarrow{\bm{\theta}}_{*}^{h_{j}}$.
  end for
Convergence rate of Full-Multiscale-SGD. To further understand the effect of the full multiscale approach, we recall the stochastic gradient descent converge rate. For the case where the learning rate converges to 00, we have that after $k$ iterations of SGD, we can bound the error of the loss by
$\left|{\mathbb{E}}\left[{\ell}({\bm{\theta}}_{k})\right]-{\mathbb{E}}\left[{\ell}({\bm{\theta}}_{*}^{h})\right]\right|\leq C_{0}{\frac{C}{k}},$ (22)
where $C$ is a constant, ${\bm{\theta}}_{*}^{h}$ is the parameter that optimizes the expectation of the loss, and $C_{0}$ is a constant that depends on the initial error at ${\bm{\theta}}_{0}$. Let ${\bm{\theta}}_{*}^{H}$ and ${\bm{\theta}}_{*}^{h}$ be the parameters that minimize the expectation of the loss for meshes with resolution $H>h$ and assume that $\|{\bm{\theta}}_{*}^{H}-{\bm{\theta}}_{*}^{h}\|\leq\gamma H$, where $\gamma$ is a constant. This assumption is justified if the loss converges to a finite value as $h\rightarrow 0$. During the Full-Multiscale-SGD iterations (Algorithm 2), we solve the problem on the coarse mesh $H$ to initialize the fine mesh $h$. Thus, after $k$ steps of SGD on the fine mesh, we can bound the error by,
$\Big{|}{\mathbb{E}}\left[{\ell}({\bm{\theta}}_{k})\right]-{\mathbb{E}}\left[{\ell}({\bm{\theta}}_{*}^{h})\right]\Big{|}\leq\gamma H{\frac{C}{k}}.$ (23)
Requiring that the error is smaller than some $\epsilon$, renders a bounded number of required iterations $k\approx\gamma C{\frac{H}{\epsilon}}$. Since in Algorithm 2, $H=2^{j}h$, the number of iterations for a fixed error $\epsilon$ is halved at each level, the iterations on the finest mesh are a fraction of those on the coarsest mesh. As shown in Table 3, optimal implementation of multiscale can speed up training by a factor of $\sim$160 compared to standard single-scale training.
5 Experimental Results
In this section, we evaluate the empirical performance of our training strategies: Multiscale-SGD (Algorithm 1) and Full-Multiscale-SGD (Algorithm 2), compared to the standard CNN training with SGD. We demonstrate the broad application of our training strategies on architectures like ResNet (He et al., 2016), UNet (Ronneberger et al., 2015), and ESPCN (Shi et al., 2016) on a wide variety of tasks ranging from image denoising, deblurring, inpainting, and super-resolution.
For each of the tasks, we compare a standard network along with their MFC versions, where we replace standard convolutions with our mesh-free convolutions (Section 3.1). We denote these modified architectures as e.g., MFC-ResNet, MFC-UNet, MFC-ESPCN, etc. Additional details on experimental settings, hyperparameters, architectures, and datasets are provided in Appendix E.
Research questions. We seek to address the following questions: 1. How effective are standard convolutions with multiscale training? 2. Do mesh-free convolutions (MFCs) perform as predicted theoretically in a multiscale setting? 3. Can multiscale training be broadly applied to typical CNN tasks, and how much computational savings does it offer compared to standard training?
Metrics. To address our questions, we focus on performance metrics (e.g., MSE or SSIM) and the computational effort for each method. As a baseline, we train the problem on a single input resolution. To be unbiased to implementation and software limitations, we define a work unit (#WU ) as one application of the model on a single image at the finest mesh (i.e., original resolution). In multiscale training, this unit decreases by a factor of 4 with each downsampling by a factor of 2. We then compare #WU  across training strategies.
Blur Estimation. We begin by solving a quadratic problem to validate the theoretical concepts discussed in Sections 2.2 and 3.2. To this end, we estimate a blurring kernel, ${{\bf G}}$, directly from data (see Nagy & Hansen (2006) for the definition). We assume data ${\bf y}$ of the form ${\bf y}={\bf G}\star{\bf u}+{\bf z}$, where ${\bf u}$ is the original (unblurred) data and ${\bf z}$ is noise. The standard MSE loss is used to estimate ${\bf G}$.
We measure computational work using #WU by running 200 iterations on the fine mesh with SGD, Multiscale-SGD, and Full-Multiscale-SGD, and early-stopping when the validation loss plateaus. Table 3 shows that using Multiscale-SGD yields significant savings, and Full-Multiscale-SGD offers even further savings. We discuss additional details for this experiment in Section D.1.
Table 3: Iterations and #WU for blur estimation. Multiscale training significantly reduces costs.
Training strategy Iterations #WU ($\downarrow$)
SGD (Single Scale) 200 $12800$
Multiscale-SGD (Ours) 200 $187$
Full-Multiscale-SGD (Ours) 145 $81$
Image Denoising. Here, one assumes data of the form ${\bf u}^{h}=t{\bf y}^{h}+(1-t){\bf z}$, where ${\bf y}^{h}$ is some image on a fine mesh $h$ and ${\bf z}\sim\mathcal{N}(0,{\bf I})$ is the noise. The noise level $t\in[0,1]$ is chosen randomly. The goal is to recover ${\bf y}^{h}$ from ${\bf u}^{h}$. The loss to be minimized is $loss({\bm{\theta}})={\frac{1}{2}}{\mathbb{E}}_{{\bf y}^{h},{\bf u}^{h},t}\|f({\bf u}^{h},t,{\bm{\theta}})-{\bf y}^{h}\|^{2}$. In Table 4, Multiscale-SGD and Full-Multiscale-SGD significantly improve training compute (as measured by #WU ) while maintaining the quality of the reconstructed denoised image on the STL10 test dataset, measured by MSE. Additional experiments on the CelebA dataset for this task are provided in Section D.2 (see, Table 5).
Table 4: Comparison of different training strategies, SGD, Multiscale-SGD and Full-Multiscale-SGD, over different standard convolution and MFC-based networks across various tasks such as denoising, deblurring, inpainting, and super-resolution.
Image Denoising, MSE ($\downarrow$)
Training strategy #WU ($\downarrow$) UNet ResNet MFC-UNet MFC-ResNet
SGD (Single-scale) 480,000 0.1918 0.1629 0.1800 0.1862
Multiscale-SGD (Ours) 74,000 0.1975 0.1653 0.1719 0.1522
Full-Multiscale-SGD (Ours) 28,750 0.1567 0.1658 0.2141 0.1744
Image Deblurring, MSE ($\downarrow$)
UNet ResNet MFC-UNet MFC-ResNet
SGD (Single-scale) 480,000 0.1510 0.1189 0.1324 0.1063
Multiscale-SGD (Ours) 74,000 0.1458 0.1211 0.1239 0.1057
Full-Multiscale-SGD (Ours) 28,750 0.1561 0.1557 0.1257 0.1191
Image Inpainting, SSIM ($\uparrow$)
UNet ResNet MFC-UNet MFC-ResNet
SGD (Single-scale) 480,000 0.8840 0.8820 0.9161 0.8745
Multiscale-SGD (Ours) 74,000 0.8627 0.8463 0.8828 0.8686
Full-Multiscale-SGD (Ours) 126,000 0.8599 0.8643 0.8749 0.8629
Image Super-resolution, SSIM ($\uparrow$)
ESPCN ResNet MFC-ESPCN MFC-ResNet
SGD (Single-scale) 30,000 0.8491 0.8366 0.8229 0.8451
Multiscale-SGD (Ours) 4,625 0.8138 0.8105 0.7988 0.8300
Full-Multiscale-SGD (Ours) 7,875 0.8042 0.8013 0.7866 0.8173
Image Deblurring. Here, one assumes data of the form ${\bf u}^{h}={\bf K}^{h}{\bf y}^{h}+{\bf z}$, where ${\bf y}^{h}$ is some image on mesh $h$, ${\bf K}^{h}$ is the blurring kernel and ${\bf z}$ is the noise. The goal is to recover ${\bf y}^{h}$ from blurred ${\bf u}^{h}$. The loss to be minimized is $loss({\bm{\theta}})={\frac{1}{2}}{\mathbb{E}}_{{\bf y}^{h},{\bf u}^{h}}\|f({\bf u}^{h},{\bm{\theta}})-{\bf y}^{h}\|^{2}$. In Table 4, our Full-Multiscale-SGD significantly accelerates training while maintaining the quality of the reconstructed deblurred image on the STL10 dataset, measured by MSE.
Image Inpainting. Here, one assumes data of the form ${\bf u}^{h}={\bf M}^{h}{\bf y}^{h}+{\bf z}$, where ${\bf y}^{h}$ is a complete image on mesh $h$, ${\bf M}^{h}$ is the image corruption operation and ${\bf z}$ is the noise. The goal is to recover ${\bf y}^{h}$ from incomplete noised data ${\bf u}^{h}$. The loss to be minimized is $loss({\bm{\theta}})={\frac{1}{2}}{\mathbb{E}}_{{\bf y}^{h},{\bf u}^{h}}\|f({\bf u}^{h},{\bm{\theta}})-{\bf y}^{h}\|^{2}$. In Table 4, our Full-Multiscale-SGD significantly accelerates training while maintaining the quality of the reconstructed inpainted image on the CelebA dataset, measured by SSIM.
Image Super-resolution. Here, we aim to predict a high-resolution image ${\bf u}^{h}$ from a low-resolution image ${\bf y}^{l}$, which is a downsampled version of ${\bf u}^{h}$. The downsampling process is modeled as ${\bf y}^{l}={\bf D}{\bf u}^{h}+{\bf z}$, where ${\bf D}(\cdot)$ is a downsampling operator (e.g., bicubic), and ${\bf z}$ is noise. The goal is to reconstruct ${\bf u}^{h}$ using a neural network $f({\bf y}^{l},{\bm{\theta}})$. The loss function is $\mathcal{L}({\bm{\theta}})=-\mathbb{E}_{{\bf y}^{l},{\bf u}^{h}}\left[\text{SSIM}\left(f({\bf y}^{l},{\bm{\theta}}),{\bf u}^{h}\right)\right].$ In Table 4, our Full-Multiscale-SGD significantly accelerates training while maintaining image quality on the Urban100 dataset, measured by SSIM.
Detailed experimental settings and visualizations for each of the above tasks have been provided in Appendix E and Appendix F, respectively. Furthermore, we provide additional experiments with deeper networks and other type of convolutions in Sections D.3 and D.4, respectively. We also compare runtimes between standard convolutions and MFCs in milliseconds and FLOPs across different image and kernel sizes and runtime per training iteration for SGD and Multiscale-SGD strategies in Section D.5.
6 Conclusion
In this paper, we introduced a novel approach to multiscale training for deep neural networks, addressing the limitations of standard mesh-based convolutions. We showed that theoretically, standard convolutions may not converge as $h\rightarrow 0$, hindering performance. To solve this, we proposed mesh-free convolutions inspired by parabolic PDEs with trainable parameters, ensuring consistent, mesh-independent convergence. Our multiscale training approach can also be coupled with standard convolutions, positioning it as an alternative to standard SGD. This approach is well-suited for high-resolution and 3D learning tasks where scalability and precision are crucial. While FFT-based MFCs are currently slower than standard convolutions, their theoretical FLOPs count is competitive, motivating future efficient implementations.
Impact Statement
This paper presents a method for improving the scalability and efficiency of CNN training through multiscale optimization. Our approach reduces training compute costs while maintaining the performance of deep networks across various tasks, which is particularly useful for high-resolution learning tasks. Additionally, our findings suggest that multiscale training can be integrated with standard convolutions across various network architectures as an alternative to conventional SGD. This paper presents work with a goal to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Background on the solution of parabolic partial differential equation
The convolution we use is defined by the solution of the parabolic PDE
$v_{t}={\cal L}v\quad\quad v(0,{\bf x})={\bf R}u$ (24)
equipped with periodic boundary conditions and on the bounded domain $[0,1]^{2}$. We now review the basic principles of parabolic PDEs to shed light on this choice.
We start with the operator ${\cal L}$ that has the form
${\cal L}v=\alpha_{xx}{\frac{\partial^{2}v}{\partial x^{2}}}+2\alpha_{xy}{\frac{\partial^{2}v}{\partial x\partial y}}+\alpha_{yy}{\frac{\partial^{2}v}{\partial y^{2}}}$ (25)
We say that the operator is elliptic if
$\alpha_{xx}\alpha_{yy}-\alpha_{xy}^{2}>0$ (26)
The ellipticity of the operator guarantees that it has an inverse (up to a constant) and that the parabolic Equation 25 is well-posed.
It is easy to verify that the eigenfunctions are given by Fourier vectors of the form
$\phi_{mn}(x,y)=\exp(2\pi i(mx+ny))$
Substituting an eigenfunction in the differential operator, we obtain that the corresponding negative eigenvalue is
$\lambda_{mn}=4\pi^{2}\left(\alpha_{xx}m^{2}+\alpha_{yy}n^{2}+2\alpha_{xy}nm\right)$
Note that the eigenvalues are negative as long as Equation 26 holds.
Consider now a time-dependent problem given by Equation 24. Consider a solution of the form
$v(t,x,y)=\sum_{mn}v_{mn}(t)\phi_{mn}(x,y)$
Substituting, we obtain an infinite set of ODEs of the form
${\frac{\partial v_{mn}}{\partial t}}=-\lambda_{mn}v_{mn}$ (27)
with the solution
$v_{mn}=\exp(-\lambda_{mn}t)v_{mn}(0).$
Since $\lambda_{mn}>0$, the solution for each component decays to 00 with faster decay for high-frequency components. Thus, the convolution is well-posed as long as the condition in Equation 26 is kept.
Appendix B Proofs
B.1 Proof to Lemma 2.1
Proof.
To prove Lemma 2.1, we further analyze Example 1 by asking: how does the gradient of the function behave as we downsample the mesh? To this end, we study the linear model in Example 1 for a single input and output channels, with a 1D convolution kernel of size 3. The loss to be optimized is defined as:
$\ell({\bm{\theta}},{\bf u}^{h},{\bf y}^{h})={\frac{1}{n}}({\bf y}^{h})^{\top}({\bm{\theta}}\star{\bf u}^{h})={\frac{1}{n}}\sum_{j}({\bm{\theta}}_{1}{\bf u}^{h}_{j-1}+{\bm{\theta}}_{2}{\bf u}_{j}^{h}+{\bm{\theta}}_{3}{\bf u}^{h}_{j+1}){\bf y}_{j}^{h},$ (28)
where ${\bm{\theta}}$ are trainable convolution weights, and ${\bf y}^{h}$ is a discretization of some function on the grid ${\bf x}^{h}$. The derivative with respect to ${\bm{\theta}}$ is given by:
${\bf g}_{1}^{h}={\frac{1}{n}}\sum{\bf y}_{j}^{h}{\bf u}_{j+1}^{h},\quad\quad{\bf g}_{2}^{h}={\frac{1}{n}}\sum{\bf y}_{j}^{h}{\bf u}_{j}^{h},\quad\quad{\bf g}_{3}^{h}={\frac{1}{n}}\sum{\bf y}_{j}^{h}{\bf u}_{j-1}^{h}.$ (29)
Let us also consider the same analysis, but on a nested mesh with mesh size $2h$, that yields:
${\bf g}_{1}^{2h}={\frac{2}{n}}\sum{\bf y}_{j}{\bf u}_{j+2},\quad\quad{\bf g}_{2}^{2h}={\frac{2}{n}}\sum{\bf y}_{j}^{h}{\bf u}_{j}^{h},\quad\quad{\bf g}_{3}^{2h}={\frac{2}{n}}\sum{\bf y}_{j}^{h}{\bf u}_{j-2}^{h}.$ (30)
If both functions $u(x)$ and $y(x)$ are smooth, then, we can use Taylor’s theorem to obtain
${\bf y}_{j}{\bf u}_{j+2}={\bf y}_{j}({\bf u}_{j+1}+h{\bf u}^{\prime}_{j+1}+{\cal O}(h^{2}))={\bf y}_{j}{\bf u}_{j+1}+{\cal O}(h),$
[…]
B.2 Proof to Lemma 3.1
Proof.
To prove the Lemma, we recall that the convolution is invariant under Fourier transform. Indeed, let $\widehat{u}(k_{x},k_{y})={\cal F}u$ and $\widehat{y}(k_{x},k_{y})={\cal F}y$, then, we can write
[…]
Since the symbol $\lim_{k_{x},k_{y}}{\cal L}(k_{x},k_{y})\rightarrow 0$ exponentially, the loss in continuous space, $\ell^{*}$ converges to a finite quantity. Moreover, the derivatives of the loss with respect to its parameters are
[…]
which also converges to a finite quantity by Parseval’s theorem.
B.3 Equivalence between regular convolutions and discretized MFCs on a fixed mesh
Lemma B.1.
Given a fixed mesh of size $h$ and a $k\times k$ convolution ${\bf K}$. There exists a MFC ${\bf K}_{{\bm{\theta}}}$ with parameters ${\bm{\theta}}$ that is discretized on the same mesh such that given $\epsilon>0$
$\|{\bf K}_{{\bm{\theta}}}{\bf u}-{\bf K}{\bf u}\|^{2}\leq\epsilon,\quad\forall{\bf u}$ (33)
The implications of the lemma is that given a standard network, it is always possible to replace the convolutions with MFC and remain within the same accuracy, as our experiments show.
Proof.
The proof for this lemma is straightforward. Notice that the convolution is made of two parts. The first is a spatial component that involves the Fourier transform and the operation in Fourier space, and the second is a $1\times 1$ convolution that combines the spatial channels. We prove the lemma for a single channel. For problems with $c$ channels, it is possible to concatenate the same structure $c$ times with different parameters.
To obtain any $k\times k$ convolution, we need to be able to express the convolution stencil by a basis of $k^{2}$ convolutions that are linearly independent. To be specific, the standard convolution can be written as
${\bf K}{\bf u}=\sum_{j=1}^{k}\sum_{i=1}^{k}{\bf K}_{ij}({\bf e}_{ij}\star{\bf u})$ (34)
where ${\bf e}_{ij}$ represents a convolution with a stencil of zeros and $1$ in the $i,j$ location and ${\bf K}_{ij}$ is the $ij$ entry of ${\bf K}$. For Mesh Free convolution, discretize on a regular mesh and using different values of $\alpha_{xy}$, then the solutions of the parabolic PDE are linearly independent (Evans, 1998). Now use any combination of the parameters ${\bm{\gamma}}$ and vary $\alpha_{xy}$ over $m\geq k^{2}$ different values. Choose $\tau$ sufficiently small such that it covers the same apparatus of the standard convolution. We obtain $k^{2}$ different spatial convolutions $\widehat{{\bf e}}_{i}$, each corresponds to a different tilting parameter $\alpha_{xy}$. Since the new set of convolutions are linearly independent, they form a basis for any $m$ convolution. In particular, it is possible to minimize the over-determined linear system for the coefficients ${\bf c}$ such that
$\min_{{\bf c}}\left\|\sum_{j=1}^{k}\sum_{i=1}^{k}{\bf K}_{ij}{\bf e}_{ij}-\sum_{i=1}^{m}{\bf c}_{i}{\widehat{\bf e}}_{i}({\bm{\theta}})\right\|^{2}$ (35)
Appendix C Computation of #WU within a multiscale framework
Definition C.1 (Working Unit (WU)).
A single working unit (WU) is defined by the computation of a model (neural network) on an input image on its original (i.e., highest) resolution.
Remark C.2.
To measure the number of working units (#WUs) required by a neural network and its training strategy, we measure how many evaluations of the highest resolution are required. That is, evaluations at lower resolutions are weighted by the corresponding downsampling factors. In what follows, we elaborate on how #WUs are measured.
We now show how to measure the computational complexity in terms of #WU for the three training strategies SGD, Multiscale-SGD, and Full-Multiscale-SGD. For the Multiscale-SGD approach, all computations happen on the finest mesh (size $h$), while for Multiscale-SGD and Full-Multiscale-SGD, the computations are performed at 4 resolutions $(h,2h,4h,8h)$. The computation of running the network on half resolution is 1/4 of the cost, and every coarsening step reduces the work by an additional factor of 4. In the Equations below, we denote use the acronyms MSGD for Multiscale-SGD, and FMSGD for Full-Multiscale-SGD From equation 2.1, we have,
${\mathbb{E}}{\bf g}^{h}({\bm{\theta}})={\mathbb{E}}\left[{\bf g}^{h}({\bm{\theta}})-{\bf g}^{2h}({\bm{\theta}})\right]+{\mathbb{E}}\left[{\bf g}^{2h}({\bm{\theta}})-{\bf g}^{4h}({\bm{\theta}})\right]+{\mathbb{E}}\left[{\bf g}^{4h}({\bm{\theta}})-{\bf g}^{8h}({\bm{\theta}})\right]+{\mathbb{E}}\left[{\bf g}^{8h}({\bm{\theta}})\right]$ (36)
With Multiscale-SGD, the number of #WU in one iteration needed to compute the ${\mathbb{E}}{\bf g}^{h}({\bm{\theta}})$ is given by,
#$\text{{\#WU\ }}_{\text{MSGD}}=N_{0}\bigg{(}1+\frac{1}{4}\bigg{)}+N_{1}\bigg{(}\frac{1}{4}+\frac{1}{16}\bigg{)}+N_{2}\bigg{(}\frac{1}{16}+\frac{1}{64}\bigg{)}+\frac{N_{3}}{64}#$ (37)
where $N_{0},N_{1},N_{2}$ and $N_{3}$ represent the batch size at different scales. With $N_{1}=2N_{0}$, $N_{2}=4N_{0}$ and $N_{3}=8N_{0}$, #WU for $I$ iterations become,
#$\text{{\#WU\ }}_{\text{MSGD}}=\frac{37N_{0}I}{16}\approx 2.31N_{0}I#$ (38)
Alternatively, seeing an equivalent amount of data, doing these same computations on the finest mesh with SGD, the #WU per iteration is given by, $N_{0}\times 1+N_{1}\times 1+N_{2}\times 1+N_{3}\times 1$ images in one iteration (where each term is computed at the finest scale). With $N_{1}=2N_{0}$, $N_{2}=4N_{0}$, and $N_{3}=8N_{0}$, the total #WU for $I$ iterations in this case, becomes,
#$\text{{\#WU\ }}_{\text{SGD}}=15N_{0}I#$ (39)
Thus, using Multigrid-SGD is roughly $6.5$ times cheaper than SGD.
[…]
#$\text{{\#WU\ }}_{\text{FMSGD}}(h)#$ $=$ $I_{h}\times\bigg{[}N_{0}^{h}\bigg{(}1+\frac{1}{4}\bigg{)}+N_{1}^{h}\bigg{(}\frac{1}{4}+\frac{1}{16}\bigg{)}+N_{2}^{h}\bigg{(}\frac{1}{16}+\frac{1}{64}\bigg{)}+\frac{N_{3}^{h}}{64}\bigg{]}$ (40)
#$\text{{\#WU\ }}_{\text{FMSGD}}(2h)#$ $=$ $\frac{I_{2h}}{4}\times\bigg{[}N_{0}^{2h}\bigg{(}1+\frac{1}{4}\bigg{)}+N_{1}^{2h}\bigg{(}\frac{1}{4}+\frac{1}{16}\bigg{)}+\frac{N_{2}^{2h}}{16}\bigg{]}$ (41)
#$\text{{\#WU\ }}_{\text{FMSGD}}(4h)#$ $=$ $\frac{I_{4h}}{16}\times\bigg{[}N_{0}^{4h}\bigg{(}1+\frac{1}{4}\bigg{)}+\frac{N_{1}^{4h}}{4}\bigg{]}$ (42)
#$\text{{\#WU\ }}_{\text{FMSGD}}(8h)#$ $=$ $\frac{I_{8h}}{64}\times N_{0}^{8h}$ (43)
where $I_{h},I_{2h},I_{4h}$ and $I_{8h}$ represent the number of training iterations at each scale and $N_{1}^{r}=2N_{0}^{r}$, $N_{2}^{r}=4N_{0}^{r}$, and $N_{3}^{r}=8N_{0}^{r}$ for each $r\in\{h,2h,4h,8h\}$ with $N_{0}^{2^{j}h}=2^{j}N_{0}$. For the denoising and deblurring tasks, we chose $I$ iterations at the coarsest scale with $I=I_{8h}=2I_{4h}=4I_{2h}=8I_{h}$. The total #WU for Full-Multiscale-SGD for these two tasks is given by,
#$\text{{\#WU\ }}_{\text{FMSGD}}=\frac{37}{16\cdot 8}N_{0}I+\frac{17}{32\cdot 4}2N_{0}I+\frac{7}{64\cdot 2}4N_{0}I+\frac{1}{64}8N_{0}I=\frac{115}{128}N_{0}I\approx 0.90N_{0}I#$ (45)
Thus, it is roughly $16$ times more effective than using SGD. For both denoising and deblurring, we chose $N_{0}=16$ and $I=2000$.
Similarly, for the inpainting and super-resolution tasks, we chose $I$ iterations at the coarsest scale with $I=I_{8h}=I_{4h}=I_{2h}=I_{h}$. We observed that a larger number of iterations per level were required for these tasks to achieve a similar accuracy as SGD. The #WU for Full-Multiscale-SGD for these three tasks is given by,
#$\text{{\#WU\ }}_{\text{FMSGD}}=\frac{37}{16}N_{0}I+\frac{17}{16}N_{0}I+\frac{7}{16}N_{0}I+\frac{1}{8}N_{0}I=\frac{63}{16}N_{0}I\approx 3.94N_{0}I#$ (46)
Thus, it is roughly $3.8$ times more effective than using SGD. For inpainting and super-resolution, we chose $N_{0}=16$ and $I=2000$, and $N_{0}=8$ and $I=250$, respectively.
Appendix D Additional results
D.1 Experimental details and training convergence analysis for the blur estimation task
(a) Standard Convolution
(b) Mesh-Free Convolution
(c) Training Loss
Figure 4: Training on a Single Scale vs. Full-Multiscale-SGD (Ours). (a) uses standard convolutions, and (b) shows Mesh-free convolutions (MFCs). Standard convolutions struggle with varying input resolutions, as seen in the gradient increase, while MFCs handle continuous training with varying resolutions. (c) plots training loss for standard and multiscale strategies with both convolution types, showing that standard convolutions are unsuitable for multiscale training, whereas MFCs enable it.
For the blur estimation task, we assume that both the image ${\bf u}$ and its blurred version ${\bf y}$ are available on a fine mesh of size $256^{2}$. The relationship is linear, given by ${\bf y}={\bf G}\star{\bf u}+{\bf z}$, where ${\bf G}$ is the blurring kernel. Given ${\bf u}$ and ${\bf y}$, we solve the optimization problem $\min_{{\bf G}}{\frac{1}{2}}{\mathbb{E}}_{{\bf u},{\bf y}}\|{\bf y}-{\bf G}\star{\bf u}\|^{2}.$
This is a quadratic stochastic optimization problem for the kernel ${\bf G}$. We simulate ${\bf u}$ as a Random Markov Field (Tenorio et al., 2011) using a random $7\times 7$ blurring kernel for ${\bf G}$. The problem is solved with standard CNNs and MFCs, comparing SGD, Multiscale-SGD, and Full-Multiscale-SGD strategies. As seen in Figure 4, for standard convolutions, the loss increases with mesh refinement, making multiscale approaches ineffective. However, our MFCs maintain stable loss across meshes, making the multiscale process efficient, with the final 50 iterations on the finest mesh.
D.2 Experiments for the denoising task on the CelebA dataset
To observe the behavior of the Full Multiscale algorithm, we performed additional experiments for the denoising task on the CelebA dataset using UNet, ResNet, MFC-UNet, and MFC-ResNet. The results have been presented in Table 5, showing that both multiscale training strategies achieved similar performance to SGD for all networks with a considerably lower number of #WU .
Table 5: Comparison of different training strategies using UNet, ResNet, MFC-UNet, and MFC-ResNet for the denoising task on the CelebA dataset.
Training strategy #WU ($\downarrow$) MSE ($\downarrow$)
UNet ResNet MFC-UNet MFC-ResNet
SGD (Single Scale) 480,000 0.0663 0.0553 0.0692 0.0699
Multiscale-SGD (Ours) 74,000 0.0721 0.0589 0.0691 0.0732
Full-Multiscale-SGD (Ours) 28,750 0.0484 0.0556 0.0554 0.0533
D.3 Experiments with deeper networks
Table 6: Comparison of different training strategies using standard convolution and MFC-based variants of ResNet18, ResNet50, and a deeper UNet for the denoising task on the STL10 dataset.
Training Strategy #WU ($\downarrow$) MSE ($\downarrow$)
ResNet18 ResNet50 MFC-ResNet18 MFC-ResNet50 UNet (5 levels) MFC-UNet (5 levels)
SGD (Single Scale) 480,000 0.1623 0.1614 0.1620 0.1611 0.1610 0.1609
Multiscale-SGD (Ours) 74,000 0.1622 0.1611 0.1617 0.1602 0.1604 0.1605
Full-Multiscale-SGD (Ours) 28,750 0.1588 0.1598 0.1598 0.1599 0.1597 0.1594
While we have shown that MFCs can replace convolutions on shallow networks as well as UNets, in our following experiment, we show that they can perform well even for deeper networks. To this end, we compare the training to ResNet-18 and ResNet-50, as well as UNets with 5 levels, for the super-resolution task on Urban 100. The results are presented in Table 6.
D.4 Comparison with Fourier Base Convolutions for the super-resolution task
Table 7: Comparison of different training strategies using ESPCN, MFC-ESPCN, and FNO-ESPCN for the super-resolution task on the Urban100 dataset.
Training strategy SSIM ($\uparrow$) ESPCN MFC-ESPCN FNO-ESPCN SGD (Single Scale) 0.841 0.842 0.744 Multiscale-SGD (Ours) 0.822 0.831 0.743 Full-Multiscale-SGD (Ours) 0.811 0.823 0.732
We conducted experiments on the Urban100 dataset to evaluate the performance of different convolutional techniques for the super-resolution task. Specifically, we compared the widely used ESPCN architecture with our proposed MFC-ESPCN (which uses parabolic, mesh-free convolutions) and FNO-ESPCN (which incorporates spectral convolutions used by the FNO paper). The results, measured using the SSIM metric, highlight that our MFC-ESPCN achieves comparable or better performance than standard convolutional methods while requiring lower computational resources. Furthermore, MFC-ESPCN outperforms the FNO-based spectral convolution approach, as it effectively captures both local and global dependencies without being limited by the truncation of frequencies in the Fourier domain. The results for the task of multi-resolution are presented in Table 7.
D.5 Comparison of runtimes between a standard convolution and MFC and between SGD and Multiscale-SGD
Figure 5: Comparison of runtimes (milliseconds) for a single standard 2D convolution with different kernel sizes ($3\times 3$, $5\times 5$, $7\times 7$, $9\times 9$ and $13\times 13$) and MFC over images of different sizes ranging from $32\times 32$ to $512\times 512$ with 3 channels. All runs were performed on the CPU.
We compared the runtimes of a single standard 2D convolution (nn.Conv2d from PyTorch) over different kernel sizes ($3\times 3$, $5\times 5$, $7\times 7$, $9\times 9$ and $13\times 13$) and image sizes ranging from $32\times 32$ to $512\times 512$ with 3 channels. Similarly, we also computed runtimes for a single MFC operation (with scaling factor $\tau=0.1$) over different image sizes. All operations were performed on an Intel(R) Xeon(R) Gold 5317 CPU @ 3.00GHz with x86_64 processor, 48 cores, and a total available RAM of 819 GB. The average of the runtimes (in milliseconds) over 500 operations for all experiments have been presented in Figure 5. In addition, we report the FLOPs required by standard convolutions with varying kernel and input image sizes, with the FLOPs required by our MFCs, in Table 8.
Table 8: Comparison of FLOPs for standard convolutions and MFCs across different image and kernel sizes. The reported FLOPs are in the format of FLOPs for standard convolutions / FLOPs for our MFCs.
Kernel size $(k\times k)\rightarrow$
Image size $(n\times n)\downarrow$
$k=32$ $k=64$ $k=128$ $k=256$
$n=5$ 2.62 ×$\times 10^{7}$ / 5.56 ×$\times 10^{7}$ 1.05 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 4.19 ×$\times 10^{8}$ / 1.23 ×$\times 10^{9}$ 1.68 ×$\times 10^{9}$ / 5.57 ×$\times 10^{9}$
$n=7$ 5.14 ×$\times 10^{7}$ / 5.56 ×$\times 10^{7}$ 2.06 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 8.22 ×$\times 10^{8}$ / 1.23 ×$\times 10^{9}$ 3.29 ×$\times 10^{9}$ / 5.57 ×$\times 10^{9}$
$n=9$ 8.49 ×$\times 10^{7}$ / 5.56 ×$\times 10^{7}$ 3.40 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 1.36 ×$\times 10^{9}$ / 1.23 ×$\times 10^{9}$ 5.44 ×$\times 10^{9}$ / 5.57 ×$\times 10^{9}$
$n=11$ 1.27 ×$\times 10^{8}$ / 5.56 ×$\times 10^{7}$ 5.08 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 2.03 ×$\times 10^{9}$ / 1.23 ×$\times 10^{9}$ 8.12 ×$\times 10^{9}$ / 5.57 ×$\times 10^{9}$
$n=13$ 1.77 ×$\times 10^{8}$ / 5.56 ×$\times 10^{7}$ 7.09 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 2.84 ×$\times 10^{9}$ / 1.23 ×$\times 10^{9}$ 1.13 ×$\times 10^{1}0$ / 5.57 ×$\times 10^{9}$
$n=15$ 2.36 ×$\times 10^{8}$ / 5.56 ×$\times 10^{7}$ 9.44 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 3.77 ×$\times 10^{9}$ / 1.23 ×$\times 10^{9}$ 1.51 ×$\times 10^{1}0$ / 5.57 ×$\times 10^{9}$
Furthermore, we compared the runtime per training iteration for both ResNet and MFC-ResNet under SGD and Multiscale-SGD training strategies. Let us consider the case of $4$ levels within a Multiscale-SGD training strategy. In this case, a multiscale computation of the gradient processes for every single image on the finest resolution, 2 images on the next resolution, $4$ on the third resolution, and $8$ on the coarsest resolution. Thus, we compare its runtimes with the time it takes to process a batch of $15$ fine-scale images with the standard approach of SGD. The results are presented in Table 9. These experiments were run on an NVIDIA A6000 GPU. We observe a significant time reduction with the multiscale training strategy. Note that the MFC-ResNet is more expensive than the standard ResNet, due to the computation of the Fourier transform, which is more expensive than a simple convolution operation.
Table 9: Comparison of runtimes (milliseconds) per training iteration using standard convolution kernels and our MFCs. The results are presented for ResNet and MFC-ResNet networks with an input image size of $256\times 256$. All runs were performed on an NVIDIA A6000 GPU.
Time (ms)
Training strategy $\rightarrow$
Network $\downarrow$
SGD
Multiscale-SGD
(Ours, 4 levels)
ResNet 21.5 1.43
MFC-ResNet 27.3 1.79
Appendix E Experimental setting
For the denoising task, the experiments were conducted on STL10 and CelebA datasets using networks, UNet (Ronneberger et al., 2015) and ResNet (He et al., 2016) with standard convolutions, and MFC-UNet and MFC-ResNet with mesh-free convolutions. For deblurring and inpainting tasks, the experiments were conducted on STL10 and CelebA datasets, respectively, using the same four networks as the denoising task. For the deblurring experiments, we used a blurring kernel $K(x,y)=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\Big{(}-\frac{x^{2}}{\sigma_{x}^{2}}-\frac{y^{2}}{\sigma_{y}^{2}}\Big{)}$ with $\sigma_{x}=\sigma_{y}=3$ to blur the input images. For the inpainting task, we introduced up to 3 (randomly chosen for each image) small rectangles with heights and widths sampled uniformly from the range $[s/12,s/6]$, where $s\times s$ is the dimension of the image. The key details of the training experimental setup for the denoising, deblurring, and inpainting tasks are summarized in Table 10. For the super-resolution task, the experiments were conducted on Urban100 dataset using networks, ESPCN (Shi et al., 2016), ResNet (He et al., 2016), MFC-ESPCN, and MFC-ResNet. The key details of the training experimental setup for the super-resolution task are summarized in Table 11.
All our experiments were conducted on an NVIDIA A6000 GPU with 48GB of memory. Upon acceptance, we will release our source code, implemented in PyTorch (Paszke et al., 2017).
Table 10: Experimental details for training for denoising, deblurring and inpainting tasks
Component Details
Dataset
STL10 and CelebA. Images from both datasets
were resized to a dimension of 64 × 64
Network architectures
UNet: 3-layer network with 32, 64, 128 filters,
and 1 residual block (res-block) per layer
ResNet: 2-layer residual network 128 hidden channels
MFC-UNet: 3-layer network with MFCs with 32
hidden channels and 1 res-block per layer
MFC-ResNet: 2-layer residual network with MFCs
with 128 hidden channels
Number of training parameters
UNet: 2,537,187
ResNet: 597,699
MFC-UNet: 2,307,411
MFC-ResNet: 532,163
Training strategies
SGD, Multiscale-SGD and
Full-Multiscale-SGD for all networks
Loss function MSE loss
Optimizer Adam (Kingma & Ba, 2014)
Learning rate $5\times 10^{-4}$ (constant)
Batch size strategy
Dynamic batch sizing is used, adjusting the batch size
upwards during different stages of training for
improved efficiency. For details, see Appendix C.
Multiscale levels 4
Iterations per level
SGD and Multiscale-SGD (all tasks): [2000, 2000, 2000, 2000],
Full-Multiscale-SGD (denoising/deblurring): [2000, 1000, 500, 250]
Full-Multiscale-SGD (inpainting): [2000, 2000, 2000, 2000]
Evaluation metrics MSE for denoising and deblurring, SSIM for inpainting
Table 11: Experimental details for training for super-resolution task
Component Details
Dataset
Urban100, consisting of paired low-resolution and high-resolution
image patches extracted for training and validation.
Network architectures
ESPCN: 5-layer super-resolution network with
64, 64, 32, 32, and 3 filters
ResNet: 9-layer ResNet-like model with 100 hidden channels per layer
MFC-ESPCN: 5-layer super-resolution network with MFCs with 32
64, 64, 32, 32, and 3 filters
MFC-ResNet: 9-layer residual network with MFCs
with 100 hidden channels per layers
Number of training parameters
ESPCN: 69,603
ResNet: 23,315
MFC-ESPCN: 80,227
MFC-ResNet: 20,755
Training strategies
SGD, Multiscale-SGD and
Full-Multiscale-SGD for all networks
Loss function Negative SSIM loss
Optimizer Adam (Kingma & Ba, 2014)
Learning rate $1\times 10^{-3}$ (constant)
Batch size strategy
Dynamic batch sizing is used, adjusting the batch size
upwards during different stages of training for
improved efficiency. For details, see Appendix C.
Multiscale levels 4
Iterations per level
SGD, Multiscale-SGD, Full-Multiscale-SGD: [250, 250, 250, 250]
Evaluation metrics SSIM
Appendix F Visualizations
In this section, we include visualization of the outputs obtained from different networks with standard convolutions, as well as our MFC-based variants for various tasks such as image denoising, deblurring, inpainting, and super-resolution. Visualizations for the denoising task on the STL-10 dataset are provided in Figures 6 and 7 and on the CelebA dataset are provided in Figures 8 and 9. Visualizations for the deblurring task on the STL-10 dataset are provided in Figures 10 and 11, and for the inpainting task on the CelebA dataset are provided in Figures 12 and 13. Visualizations for the super-resolution task on the Urban100 dataset are provided in Figures 14 and 15.
Appendix G Comparing feature maps between CNNs and MFCs
As seen in the experimental results, standard convolutions can benefit from the multiscale framework even though they may not converge in theory. We hypothesize that this is due to the smoothing properties of the filters that are obtained in the training of CNNs. To this end, we use the denoising experiment and plot the feature maps after the first layer in a ResNet-18. A few of the feature maps for our convolutions, as well as regular convolutions, are presented in Figure 16. As can be seen in the figure, both our filters and standard CNNs tend to smooth the noisy data and generate smoother feature maps. This can account for the success of standard CNN training in this experiment.
Figure 6: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the denoising task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 7: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the denoising task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 8: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the CelebA dataset for the denoising task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 9: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the CelebA dataset for the denoising task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 10: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the deblurring task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 11: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the deblurring task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 12: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the inpainting task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 13: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the inpainting task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 14: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the Urban100 dataset for the super-resolution task. The first column displays the low-resolution data (same for all rows), followed by results from ESPCN, ResNet, MFC-ESPCN, and MFC-ResNet.
Figure 15: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the Urban100 dataset for the super-resolution task. The first two columns display the low-resolution data and original image (same for all rows), followed by results from ESPCN, ResNet, MFC-ESPCN, and MFC-ResNet.
Figure 16: First row: Input images. Second row: feature maps obtained after the first layer of a ResNet-18 with our MFCs. Third row: feature maps obtained after the first layer of a ResNet-18 with standard convolutions. Both networks are trained on the denoising task on the STL10 dataset.